A fireman of mass 80 kg slides down a pole during a rescue mission . The force of friction is constant at 720 N .Find the acceleration of the man.(g=10 m per second square )..

 fireman of mass 80 kg slides down a pole during a rescue mission . The   force of friction is constant at 720 N .Find the acceleration of the man.(g=10 m per second square )..

Grade:

1 Answers

Arun
5 years ago
Dear Deepak
 
Resultant force = 800 – 720 = 80N
 
now 
 
80 = m * a
 
80 = 80 * a
 
a = 1 m/sec^2
 
Regards
Arun (askIITians forum expert)